OpenStudy (anonymous):

how to solve using the quadratic formula 4m(squared) - 25 = 0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

but I don't see b?

OpenStudy (anonymous):

a=4 c=25 .. what is b? 1?

hartnn (hartnn):

so b=0. a=4 is correct but c= -25 don't forget *minus* sign.

OpenStudy (anonymous):

right right, ok, so plug in b as 0 in the calc

hartnn (hartnn):

yup.

OpenStudy (anonymous):

i love you!

hartnn (hartnn):

i love u too, dear :)

hartnn (hartnn):

so did u get +5/2 or -5/2 as your answer ?

OpenStudy (anonymous):

no :( I got 40 somehow.. i don't think i'm putting it in right

OpenStudy (anonymous):

can you help me work it out? I'm not sure how this one works...

hartnn (hartnn):

sure. \(\huge{\frac{-b \pm \sqrt{b^2-4ac}}{2a}}=\huge{\frac{-0 \pm \sqrt{^2-4(4)(-25)}}{2a}}\) what is -4 * (4) *(-25)

OpenStudy (anonymous):

-4 * (4) = - 16 * -25 = 400

hartnn (hartnn):

and square root of 400 ?

OpenStudy (anonymous):

20

hartnn (hartnn):

yup, but u have + or minus sign also so numerator is +20 or -20 denominator = ?

OpenStudy (anonymous):

8

hartnn (hartnn):

so what is +20/8 or -20/8 ?

OpenStudy (anonymous):

5/2 or -5/2

OpenStudy (anonymous):

aaaaahhhhhhhhhhhh

hartnn (hartnn):

thats your final answer. any doubts anywhere.

OpenStudy (anonymous):

how do i display this as the answer ?

hartnn (hartnn):

i don't know how your system accepts. possible representations are 5/2,-5/2 2.5,-2.5

OpenStudy (anonymous):

thank you so much for your help!!

hartnn (hartnn):

my pleasure most welcome :)